OpenStudy (anonymous):

he sum of two consecutive odd integers is 28. What are the integers? Which let statement(s) would you use?

OpenStudy (anonymous):

If you divide 28 by 2. You get 14. Since we need it odd, you can have 13 and 15. Since you subtract from 1 - 14 (13), you have to add to the other 1 + 14 (15) 13 + 15 = ?

OpenStudy (anonymous):

the statement is "The sum of two consecutive odd integers is 28" please use the info ive given you in previous questions to figure out the next problems by yourself.

OpenStudy (anonymous):

Let the two consecutive odd integers be x and (x+2) thus according to question x+x+2= 28 2x=28-2 2x=26 x=26/2 x=13 hence x+2= 13+2 =15 So 13 and 15 are the required integers. @tica

OpenStudy (anonymous):

@dpasingh @QuestionsWolf it asks what statements to use to figure out the question ._.

OpenStudy (anonymous):

i did. this question bout to b closed thmc everyone.

OpenStudy (anonymous):

@BeckieBird Statement is given below: Let the two consecutive odd integers be x and (x+2)

OpenStudy (anonymous):

thnx.

OpenStudy (anonymous):

oh ok